# Distributed Inter-Strand Coupling Current Model for Finite Element Simulations of Rutherford Cables

### v1

@@ 5.2.1 Applied field along $\hat{\boldsymbol{e}}_{y}$
The influence of the spatial discretization on $Q_{\text{IS}}$ is illustrated in Fig. 13. From mesh $2$ to mesh $5$, the relative difference between calculated peak losses is less than $1.8\%$. This indicates that a relatively coarse mesh can be chosen in practice, e.g., mesh 2 or 3. The difference between the results increases with frequency, because the field distributions become increasingly non-uniform and concentrated at the edges of the cable when the frequency increases (see Fig. 11).
@@ 5.2.2 Effect of the field angle
The evolution of $Q_{\text{IS}}$ when the field angle $\theta$ varies is shown in Fig. 14. As the field angle decreases, the contribution from crossing IS coupling currents (associated with the peak around $2$ Hz) progressively decreases, and adjacent IS coupling currents become dominant, with a significantly higher peak loss frequency (around $2$ kHz). Local field maps in the case $\theta=30\degree$ are illustrated in Fig. 15 at three frequencies. The local field maps show that the homogenized cable model reproduces faithfully the cable response.
@@ 5.2.3 Transport current and applied field
The homogenized model can also be used with a transport current excitation without any further parameter tuning. The current density distribution and magnetic flux density lines obtained for a harmonic transport current at $f=10$ Hz and $f=100$ Hz, and no background field, are shown in Fig. 16. As the frequency increases, the current density evolves from an almost uniform to a non-uniform distribution, with field concentration at the edges of the cable. The associated loss per cycle and per unit length $Q_{\text{IS}}$ is shown in Fig. 17. Any combination of transport current and external field can be simulated. The agreement between the homogenized model and the reference model is very good in all cases, with a maximum error of $3.5\%$ for $f\lesssim 500$ Hz.
@@
Figure 17: $Q_{\text{IS}}$ as a function of the frequency in case of (i) a harmonic external field of amplitude $10$ mT along $\hat{\boldsymbol{e}}_{y}$ ($\theta=90\degree$), (ii) a harmonic transport current of amplitude $1000$ A, or (iii) a combination of both, in phase. Note that because the problem is linear, the loss in case (iii) is the sum of the loss in cases (i) and (ii). Dashed curves are from the reference model and solid curves are from the homogenized model.
@@ 5.3.1 Effect of the transposition length
Figure 18 illustrates that the DISCC model parameter $\bar{\boldsymbol{\sigma}}_{\text{IS}}$, Eq. (15), allows to correctly capture the influence of the transposition length $p_{\text{s}}$ on the response of the cable.
@@ 5.3.2 Effect of the number of strands
The effect of the number of strands $N_{\text{s}}$ on $Q_{\text{IS}}$ is also correctly captured by the DISCC model via $\bar{\boldsymbol{\sigma}}_{\text{IS}}$, Eq. (15). This is shown in Fig. 19, where $N_{\text{s}}$ is varied while keeping the strand width $W_{\text{s}}$ constant (hence the cable width $W_{\text{c}}=W_{\text{s}}N_{\text{s}}/2$ varies) and the periodicity length $\ell_{\text{s}}$ constant (hence the transposition length $p_{\text{s}}=\ell_{\text{s}}N_{\text{s}}$ varies).
@@ 5.3.3 Effect of the $r_{\text{c}}/r_{\text{a}}$ ratio
Finally, the effect of the $r_{\text{c}}/r_{\text{a}}$ ratio on the cable response is illustrated in Fig. 20, where $r_{\text{a}}$ is kept constant and equal to $2.06\times 10^{-11}$ $\Omega$ m2 while $r_{\text{c}}$ is varied. This illustrates the influence of a resistive strip, or core, which is sometimes introduced between the top and bottom layers of strands, as depicted in Fig. 3, in order to increase $r_{\text{c}}$ [32].
The homogenized model captures the influence of the $r_{\text{c}}/r_{\text{a}}$ ratio to some extent. The peak loss value is still accurately reproduced (relative difference of $1.5\%$ for $r_{\text{c}}/r_{\text{a}}=25$), but as the ratio $r_{\text{c}}/r_{\text{a}}$ increases, the difference between the reference and homogenized model curves increases after the peak of $Q_{\text{IS}}$. This is due to the adjacent IS coupling currents becoming the dominant coupling contribution when $r_{\text{c}}/r_{\text{a}}\gg 1$, while the DISCC model equations are primarily based on the crossing IS coupling current dynamics for a field along $\hat{\boldsymbol{e}}_{y}$, as mentioned in Section 4.1.1.
@@ 6 Verification of the nonlinear model
In the previous section, the DISCC model was verified in a linear setting. In this section, we extend the analysis to a nonlinear setting, comparing results of the reference and homogenized cable models including all the loss and magnetization mechanisms as decribed in Sections 3.2 and 4.2, respectively. Both models are nonlinear, and simulations are performed in the time domain. We consider the same flat cable as in the previous section, described by properties in Table 1. Inter-strand contact resistances are still assumed to be uniform over the cable, with values given in Table 1.
We describe the models for the internal strand dynamics in Sections 6.1 and 6.2, for the reference and homogenized cable models, respectively. We then verify the homogenized cable model results in the case of a uniform magnetic field excitation in Section 6.3, and a non-uniform magnetic field excitation in Section 6.4. Finally, we discuss and verify the differential inductance in Section 6.5.
@@ 6.1 Reference cable model parameters
The relationship between the strand voltage (per unit length) and the strand current $I_{i}$, either due to a net imposed transport current in the cable, or to a net strand current associated with IS coupling currents, is described by a combination of a ROHF model [27] of $M=5$ cells, with parameter values given in Table 3 in B.2, and a current sharing relationship defined in B.3 which describes a smooth transition between the power law and a strand resistance in the normal state.
@@ 6.2 Homogenized cable model parameters
The inclusion of strand dynamics in the homogenized cable model is done as described in Section 4.2. The ROHM model is used as a local magnetic constitutive law, but compared to the reference cable model, the mixture of strands and gaps is homogenized. With the strand filling factor $a_{\text{s}}=0.862$, we therefore define:
@@
Compared to the reference cable model, the relationship between voltage per unit length (or electric field) and strand transport current is replaced by an electric constitutive law written in terms of the homogenized current density. The same ROHF model can be used, but with scaled coercivity values $\tilde{\xi}_{k}=a_{\text{s}}\,\xi_{k}/A_{\text{s}}$, $\forall k\in\{1,\dots,M\}$. The current sharing relationship is adapted similarly, as described in B.3.
@@ 6.3 Verification for uniform magnetic field excitation
A magnetic field $\boldsymbol{h}(t)=h_{\text{max}}\sin(2\pi ft)\hat{\boldsymbol{e}}_{y}$ of various amplitudes $h_{\text{max}}$ and frequencies $f$ is applied along $\hat{\boldsymbol{e}}_{y}$. Starting from a virgin state, the simulation is performed from $t=0$ to $t=1.5/f$, using $120$ equidistant time steps. Iterations are performed with a Newton-Raphson iterative scheme, using the analytical Jacobian matrices and a convergence criterion of $10^{-6}$ (-) on the relative change of the instantaneous power between two successive iterations. Convergence was obtained in every simulation for all the time steps.
@@
Figure 21: Energy loss per cycle for an applied field along $\hat{\boldsymbol{e}}_{y}$ with three different amplitudes as a function of frequency. Solid curves are results from the homogenized cable model. Dashed curves are results from the reference cable model.
Results are shown in Fig. 21 as a function of frequency (interpolated from $25$ frequency values) for $4$ magnetic field amplitudes. The agreement between both models on the total loss $Q_{\text{tot}}$ is very good, with a relative difference smaller than $10\%$ for $f<5$ kHz. It is worth emphasizing that, apart from scaling the ROHM, ROHF and current sharing models by the strand filling factor, no extra parameter tuning was necessary to include the full strand dynamics in the model obtained in Section 5.
@@
The rate-dependency of the curve $Q_{\text{hyst}}$ is due to a combination between uncoupled filament magnetization, coupled filament magnetization, transport current hysteresis, and shielding by eddy currents at high frequencies. In particular, the second peak in the $Q_{\text{hyst}}$ curve for $\mu_{0}h_{\text{max}}=0.2$ T around $500$ Hz is due to the ROHF model contribution, i.e., the hysteresis (and eddy) loss due to strand transport currents flowing as a result of the IS coupling.
Note that the agreement between the $Q_{\text{IS}}$ curves of the reference and homogenized models at high frequencies is better compared to the linear case, e.g., Fig. 18.
@@ 6.4 Verification for non-uniform magnetic field excitation
To verify the homogenized model under a non-uniform excitation, we consider the same cable as above, but subject to the magnetic field generated by an excitation coil placed in its vicinity, as represented in Fig. 22. The coil is made of two parts of height $1.4$ mm and width $5$ mm, separated by a $4$ mm gap. The gap between the coil and the cable is $1$ mm. Each half of the coil carries a total transport current $I_{\text{coil}}(t)$ flowing in opposite direction (modelled as stranded conductors [36]). Such a configuration is relevant for the study of novel quench protection techniques such as E-CLIQ [6].
@@
$I_{\text{t}}(t)=\left\{\begin{aligned} &0,\quad\ t\leq-2~\text{s},\\ &I_{\text{t,max}}(t+2),\quad\ -2~\text{s}<t\leq-1~\text{s}\\ &I_{\text{t,max}},\quad\ t>-1~\text{s}.\end{aligned}\right$. (28)
The total coil current $I_{\text{coil}}(t)$ is defined as
@@
Despite the strongly non-uniform field distribution, the agreement between both models is excellent in all cases. Field distributions in Fig. 25 illustrate how the homogenized model reproduces the reference solution with distributed vector fields, also for the magnetization vector $\mu_{0}\boldsymbol{m}=\boldsymbol{b}-\mu_{0}\boldsymbol{h}$ described by the ROHM model.
Note that the staircase shape of the hysteresis loss curve (and hence also of the total loss curve) is due to the description of the magnetization and internal flux response using a discrete number of cells $N$ and $M$, with the ROHM and ROHF models [16], respectively, see Fig. 29. These cells are triggered one by one as the magnetic field or transport current varies. If wanted, this staircase effect can be reduced by increasing $N$ and $M$.
6.5 Verification of the differential inductance
In addition to causing loss and field distortions, IS coupling currents and internal strand dynamics affect the differential inductance of Rutherford cables [3, 4]. This leads to faster magnet discharges than what would be obtained with an ideal inductor and this effect is therefore important to quantify and evaluate for the design of quench protection systems. In this section, we verify the differential inductance calculated by the homogenized model.
We define the differential inductance per unit length $L_{\text{d}}$ (H/m) as follows:
@@ 6.5 Verification of the differential inductance
We consider the same cable as before. No external field is applied and the transport current is imposed. From a precomputed steady-state, $I_{\text{t}}$ is ramped down linearly from $10$ kA at $t=0$ s to 0 A at $t=0.5$ s. The differential inductances calculated by both models are shown in Fig. 26, including different types of dynamics in the models.
@@
In case (ii), we use the IS contact resistance values of Table 1. The differential inductances computed by both models match very well and present an initial transient with lower differential inductance, with a characteristic decay time of approximately $12$ ms, which is therefore the natural time constant for the IS dynamics related to transport current in this cable. Note that this time constant is consistent with the peak frequency $f_{\text{c}}\approx 13$ Hz observed in the loss curve of Fig. 17 in the case with transport current only. Indeed, we have $1/(2\pi f_{\text{c}})\approx 12$ ms.
@@
Finally, case (iv) includes all the internal strand dynamics by enabling all components of the ROHM, ROHF, and current sharing models. The effect of filament hysteresis in the strands decreases the differential inductance even further. This adds a very long time constant that is due to the induced current in the strands that flow predominantly in the superconductor. The agreement between the reference and homogenized models is very satisfying for both cases (iii) and (iv).
@@ 7 Stack of cables and $h$-$\phi$-$a$-formulation
In some cases, however, the $h$-$\phi$-formulation is not the most convenient nor the optimal choice. For example, in presence of ferromagnetic materials, other formulations such as the $a$-formulation offer better convergence properties [39]. In that case, in order to keep the DISCC model equations unchanged, the mixed $h$-$\phi$-$a$-formulation can be used [39, 45, 53]. In the $h$-$\phi$-$a$-formulation, one part of the domain, $\Omega_{a}$, is solved by the $a$-formulation in terms of the magnetic vector potential, $\boldsymbol{a}$ (with $\textrm{{curl}}\,\boldsymbol{a}=\boldsymbol{b}$), whereas the complementary part, $\Omega_{h}$, is solved with the $h$-$\phi$-formulation. Here, a simple choice consists in putting the cables in $\Omega_{h}$, and all remaining regions in $\Omega_{a}$. The common boundaries between $\Omega_{a}$ and $\Omega_{h}$ are the cable boundaries $\Gamma_{\text{c}}$. In order to avoid stability issues on these boundaries, second-order basis functions must be introduced on $\Gamma_{\text{c}}$ for either $\boldsymbol{a}$ or $\boldsymbol{h}$ [53]. Enriching $\boldsymbol{a}$ is simpler, in order not to interfere with the DISCC model inside the cables. A local enrichment with hierarchical basis functions supported only by edges on $\Gamma_{\text{c}}$ is sufficient [53].
@@
The results obtained with both formulations are compared in Fig. 28. The calculated power loss values are almost identical, with a relative difference smaller than 11%, which is due to the coarse spatial discretization. Field distributions obtained with the $h$-$\phi$-$a$-formulation are shown in Fig. 27 (solutions of the $h$-$\phi$-formulation are not shown but are visually the same). For this example without ferromagnetic material, the computational performance of both formulations is similar.

### v2

@@ 5.2.2 Effect of the field angle
The evolution of $Q_{\text{IS}}$ when the field angle $\theta$ varies is shown in Fig. 14. As the field angle decreases, the contribution from crossing IS coupling currents (associated with the peak around $2$ Hz) progressively decreases, and adjacent IS coupling currents become dominant. Adjacent IS coupling currents are characterized by a significantly faster dynamics, with a peak loss per cycle around $2$ kHz. The maximum difference between models for $f\lesssim 500$ Hz occurs for $\theta=0\degree$ and is equal to $11.6\%$. Local field maps in the case $\theta=30\degree$ are illustrated in Fig. 15 at three frequencies. The local field maps show that the homogenized cable model reproduces faithfully the cable response, with a combination of crossing and adjacent IS coupling currents.
@@ 5.2.3 Transport current and applied field
The homogenized model can also be used with a transport current excitation without any further parameter tuning. The current density distribution and magnetic flux density lines obtained for a harmonic transport current at $f=10$ Hz and $f=100$ Hz, and no background field, are shown in Fig. 16. As the frequency increases, the current density evolves progressively from an almost uniform to a non-uniform distribution, with field concentration at the edges of the cable. The associated loss per cycle and per unit length $Q_{\text{IS}}$ is shown in Fig. 17.
The evolution of the loss per cycle is similar to that of the applied field case. The loss per cycle first scales proportional to $f$ at low frequencies, then reaches a peak when the IS coupling current amplitude becomes comparable with the transport current, and finally decreases at higher frequencies, at a varying rate that depends on the shape of the cable. As before, the prediction of both model worsens for $f\gtrsim 500$ Hz, because of intrinsic differences between the models.
Any combination of transport current and external field can be simulated. The agreement between the homogenized model and the reference model is very good in all cases, with a maximum error of $3.5\%$ for $f\lesssim 500$ Hz.
@@
Figure 17: $Q_{\text{IS}}$ as a function of the frequency in case of (i) a harmonic external field of amplitude $10$ mT along $\hat{\boldsymbol{e}}_{y}$ ($\theta=90\degree$), (ii) a harmonic transport current of amplitude $1000$ A, or (iii) a combination of both, in phase. Dashed curves are from the reference model. Solid curves are from the homogenized model.
@@ 5.3.1 Effect of the transposition length
Figure 18 illustrates that the DISCC model parameter $\bar{\boldsymbol{\sigma}}_{\text{IS}}$, Eq. (15), correctly captures the influence of the transposition length $p_{\text{s}}$ on the response of the cable. Varying the transposition length leads to a horizontal shift of the $Q_{\text{IS}}$ curve, whose position is inversely proportional to $p_{\text{s}}^{2}$.
@@ 5.3.2 Effect of the number of strands
The effect of the number of strands $N_{\text{s}}$ on $Q_{\text{IS}}$ is also correctly captured by the DISCC model via $\bar{\boldsymbol{\sigma}}_{\text{IS}}$, Eq. (15). This is shown in Fig. 19, where $N_{\text{s}}$ is varied while keeping the strand width $W_{\text{s}}$ constant (the cable width $W_{\text{c}}=W_{\text{s}}N_{\text{s}}/2$ varies) and the periodicity length $\ell_{\text{s}}$ constant (the transposition length $p_{\text{s}}=\ell_{\text{s}}N_{\text{s}}$ varies). The results align very well with those of the reference model, even for a low number of strands.
@@ 5.3.3 Effect of the $r_{\text{c}}/r_{\text{a}}$ ratio
Finally, the effect of the $r_{\text{c}}/r_{\text{a}}$ ratio on the cable response is illustrated in Fig. 20, where $r_{\text{a}}$ is kept constant and equal to $2.06\times 10^{-11}$ $\Omega$ m2 while $r_{\text{c}}$ is varied. This illustrates the influence of a resistive strip, or core, which is sometimes introduced between the top and bottom layers of strands, as depicted in Fig. 3, in order to increase $r_{\text{c}}$, or better control its value [37].
The homogenized model captures the influence of the $r_{\text{c}}/r_{\text{a}}$ ratio to some extent. The peak loss value is still accurately reproduced (relative difference of $1.5\%$ for $r_{\text{c}}/r_{\text{a}}=25$), but as the ratio $r_{\text{c}}/r_{\text{a}}$ increases, the difference between the reference and homogenized model curves increases after the peak of $Q_{\text{IS}}$. This is due to the adjacent IS coupling currents becoming the dominant coupling contribution when $r_{\text{c}}/r_{\text{a}}\gg 1$, while the DISCC model equations are primarily based on the crossing IS coupling current dynamics for a field along $\hat{\boldsymbol{e}}_{y}$, as was mentioned in Section 4.1.1.
@@ 6 Verification of the nonlinear model
In the previous section, the DISCC model was verified in detail by considering only the IS coupling current dynamics. In this section, we extend the analysis by including all the remaining loss and magnetization mechanisms. We consider models presented in Sections 3.2 and 4.2, respectively. Both models are nonlinear, and simulations are performed in the time domain. We consider the same flat cable as in the previous section, described by properties in Table 1. IS contact resistances are still assumed to be uniform, with values given in Table 1.
We describe the models for the internal strand dynamics in Sections 6.1 and 6.2, for the reference and homogenized cable models, respectively. We then verify the homogenized cable model results in the case of a uniform magnetic field excitation in Section 6.3, and a non-uniform magnetic field excitation in Section 6.4. Finally, we discuss and verify the calculated inductance in Section 6.5.
@@ 6.1 Reference cable model parameters
The relationship between the strand voltage (per unit length) and the strand current $I_{i}$, either due to a net imposed transport current in the cable, or to a net strand current associated with IS coupling currents, is described by a combination of a ROHF model [20] using a chain of $M=5$ cells with parameter values given in Table 3 in A.2, and the current sharing relationship defined in A.3 which describes a smooth transition between the power law and a strand resistance in the normal state.
@@ 6.2 Homogenized cable model parameters
The inclusion of strand dynamics in the homogenized cable model is done as described in Section 4.2. The ROHM model is used as a local magnetic constitutive law, but compared to the reference cable model, the mixture of strands and gaps is homogenized. With the strand filling factor $a_{\text{s}}=A_{\text{c}}/(W_{\text{c}}H_{\text{c}})=0.862$, we therefore define:
@@
Compared to the reference cable model, the relationship between voltage per unit length (or electric field) and strand transport current is replaced by an electric constitutive law written in terms of the homogenized current density. The same ROHF model can be used, but with scaled coercivity values $\tilde{\xi}_{k}=a_{\text{s}}\,\xi_{k}/A_{\text{s}}$, $\forall k\in\{1,\dots,M\}$. The current sharing relationship is adapted similarly, as described in A.3.
@@ 6.3 Verification for uniform magnetic field excitation
A magnetic field $\boldsymbol{h}_{\text{app}}(t)=h_{\text{app}}\sin(2\pi ft)\hat{\boldsymbol{e}}_{y}$ of various amplitudes $h_{\text{app}}$ and frequencies $f$ is applied along $\hat{\boldsymbol{e}}_{y}$. Starting from a virgin state, the simulation is performed from $t=0$ to $t=1.5/f$, using $120$ equidistant time steps. Iterations are performed with a Newton-Raphson iterative scheme, using the analytical Jacobian matrices and a convergence criterion of $10^{-6}$ (-) on the relative change of the instantaneous power between two successive iterations. Convergence was obtained in every simulation for all the time steps.
@@
Figure 21: Estimate of the energy loss per cycle for an applied field along $\hat{\boldsymbol{e}}_{y}$ ($\theta=90\degree$) with three different amplitudes as a function of frequency. Solid curves are results from the homogenized cable model. Dashed curves are results from the reference cable model.
Results are shown in Fig. 21 as a function of frequency (interpolated from $25$ frequency values) for $4$ magnetic field amplitudes. The agreement between both models on the total loss $Q_{\text{tot}}$ is very good up to high frequencies, with a relative difference smaller than $10\%$ for $f<5$ kHz. It is worth emphasizing that, apart from scaling the ROHM, ROHF and current sharing models by the strand filling factor, no extra parameter tuning was necessary to include the full strand dynamics in the model obtained in Section 5.
@@
The rate-dependency of the curve $Q_{\text{hyst}}$ is due to a combination between uncoupled filament magnetization, coupled filament magnetization, transport current hysteresis, and shielding by eddy currents at high frequencies. In particular, the second peak in the $Q_{\text{hyst}}$ curve for $\mu_{0}h_{\text{app}}=0.2$ T around $500$ Hz is due to the ROHF model contribution, i.e., the hysteresis (and eddy) loss due to strand transport currents flowing as a result of the IS coupling.
@@ 6.4 Verification for non-uniform magnetic field excitation
To verify the homogenized model under a non-uniform excitation, we consider the same cable as above, but subject to the magnetic field generated by an excitation coil placed in its vicinity, as represented in Fig. 22. The coil is made of two parts of height $1.4$ mm and width $5$ mm, separated by a $4$ mm gap. The gap between the coil and the cable is $1$ mm. Both halves of the coil carry the same total transport current $I_{\text{coil}}(t)$, distributed uniformly, but flowing in opposite directions (modelled as stranded conductors [41]). Such a configuration is relevant for the study of the novel quench protection technique E-CLIQ [9]. This technique relies on inducing power loss in superconducting cables by generating fast field transients with an external coil, with the objective to initiate quench quickly.
@@
$I_{\text{t}}(t)=\left\{\begin{aligned} &0,\quad\ t\leq-2~\text{s},\\ &I_{\text{t,max}}(t+2),\quad\ -2~\text{s}<t\leq-1~\text{s}\\ &I_{\text{t,max}},\quad\ t>-1~\text{s},\end{aligned}\right$. (28)
see Fig. 23. The total coil current $I_{\text{coil}}(t)$ is defined as
@@
Figure 23: Transport current and current in the external coil.
@@
Despite the strongly non-uniform field distribution, the agreement between both models is very good in all cases. Depending on the frequency, distinct loss mechanisms dominate. Field distributions in Fig. 26 illustrate how the homogenized model reproduces the reference solution with distributed vector fields, also for the magnetization vector $\mu_{0}\boldsymbol{m}=\boldsymbol{b}-\mu_{0}\boldsymbol{h}$ described by the ROHM model.
Note that the staircase shape of the hysteresis loss curve (and hence also of the total loss curve) is due to the description of the magnetization and internal flux response using a low number of cells $N$ and $M$, with the ROHM and ROHF models [19], respectively, see Fig. 30. These cells are triggered one by one as the magnetic field or transport current varies. If wanted, this staircase effect can be reduced by increasing $N$ and $M$.
6.5 Verification of the inductance
In addition to causing loss and field distortions, IS coupling currents and internal strand dynamics affect the inductance of Rutherford cables [3, 4]. This leads to faster magnet discharges than what would be obtained with an ideal inductor and this effect is therefore important to quantify for the design of quench protection systems. In this section, we verify the inductance calculated by the homogenized model.
As a measure of the instantaneous apparent inductance, we consider the differential inductance per unit length $L_{\text{d}}$ (H/m) defined as follows:
@@ 6.5 Verification of the inductance
We consider the same cable as before. No external field is applied and the transport current is imposed. From a precomputed steady-state, $I_{\text{t}}$ is ramped down linearly from $10$ kA at $t=0$ s to 0 A at $t=0.5$ s. The differential inductances calculated by both models are shown in Fig. 27 as a function of time, including different types of dynamics in the models.
@@
In case (ii), we use the IS contact resistance values of Table 1. The differential inductances computed by both models match very well and present an initial transient with lower differential inductance, with a characteristic decay time of approximately $12$ ms, which is a natural time constant for the IS dynamics related to transport current in this cable. Note that this time constant is consistent with the peak frequency $f_{\text{c}}\approx 13$ Hz observed in the loss curve of Fig. 17 in the case with transport current only. Indeed, we have $1/(2\pi f_{\text{c}})\approx 12$ ms.
@@
Finally, case (iv) includes all the internal strand dynamics by enabling all components of the ROHM, ROHF, and current sharing models. The effect of filament hysteresis in the strands decreases the differential inductance even further. This adds a very long time constant that is due to the induced current in the strands that flow predominantly in the superconductor. The agreement between the reference and homogenized models is very satisfying in all cases, with a maximum relative difference smaller than $0.2\%$ for case (iv).
@@ 7 Stack of cables and $h$-$\phi$-$a$-formulation
In some cases, however, the $h$-$\phi$-formulation is not the most convenient nor the optimal choice. For example, in presence of ferromagnetic materials, other formulations such as the $a$-formulation offer better convergence properties [47]. In that case, in order to keep the DISCC model equations unchanged, the mixed $h$-$\phi$-$a$-formulation can be used [47, 46, 55]. In the $h$-$\phi$-$a$-formulation, one part of the domain, $\Omega_{a}$, is solved by the $a$-formulation in terms of the magnetic vector potential, $\boldsymbol{a}$ (with $\textrm{{curl}}\,\boldsymbol{a}=\boldsymbol{b}$), whereas the complementary part, $\Omega_{h}$, is solved with the $h$-$\phi$-formulation. Here, a simple choice consists in putting the cables in $\Omega_{h}$, and all remaining regions in $\Omega_{a}$. The common boundaries between $\Omega_{a}$ and $\Omega_{h}$ are the cable boundaries $\Gamma_{\text{c}}$. In order to avoid numerical stability issues on these boundaries (spurious oscillations), it was shown in [55] that second-order basis functions must be introduced on $\Gamma_{\text{c}}$ for either $\boldsymbol{a}$ or $\boldsymbol{h}$. Enriching $\boldsymbol{a}$ is simpler, in order not to interfere with the DISCC model inside the cables. A local enrichment with hierarchical basis functions supported only by edges on $\Gamma_{\text{c}}$ is sufficient [55].
@@
The results obtained with both formulations are compared in Fig. 29. The calculated power loss values are almost identical, with a relative difference smaller than 11%, which is due to the relatively coarse spatial discretization. Field distributions obtained with the $h$-$\phi$-$a$-formulation are shown in Fig. 28 (solutions of the $h$-$\phi$-formulation are not shown but are visually the same). For this example without ferromagnetic material, the computational performance of both formulations is similar.